OpenStudy (jenniferjuice):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21. −45 0 2 63

hartnn (hartnn):

you just need to solve for x in 6(x+2) − 3 = 21 try ?

OpenStudy (loser66):

@hartnn have to find inverse function first.

hartnn (hartnn):

thats another way ...what i posted is a shortcut

hartnn (hartnn):

conventional way is to find the inverse function and then plug in x=21 with 6(x+2) − 3 = 21 you'll get the same answer

OpenStudy (loser66):

Wow.!!! very clever!! It's new to me but it's perfectly good. Thanks @hartnn

hartnn (hartnn):

welcome ^_^ lets see whether @jenniferjuice could get the value of x.

OpenStudy (jenniferjuice):

i dont get it do i plug in 21 for x?

hartnn (hartnn):

just solve for x in 6(x+2) − 3 = 21 start by adding 3 on both sides

OpenStudy (jenniferjuice):

x=2 ?

hartnn (hartnn):

\(\huge \checkmark \)